# Generated Contents Enrichment

### v2

@@ III-C Stage 1: Generative Adversarial SG Enrichment
Our Scene Graph Enricher $G_{sg}$ processes the input graph $g$ as:
$\hat{g}=G_{sg}(g).$ (6)
This operation yields an enriched graph $\hat{g}$. Afterward, $\hat{g}$ is reintroduced into the $G_{sg}$ process, initiating an iterative procedure for further enrichment. This recursive approach allows for the scene’s gradual improvement and refinement.
$G_{sg,obj}$ is modeled with the proposed GCN. Additional layers and skip connections are introduced to deepen the GCN, avoiding issues like vanishing gradients. $G_{sg,obj}$ results in the creation of an enriching object denoted as $\hat{o}$, where
$\hat{o}=G_{sg,obj}(g).$ (7)
This newly generated enriching object is then added to the original graph $g$. The augmented graph, consisting of both the original objects and the newly created object, continues to serve as the input for the succeeding steps of the enrichment process.
$G_{sg,edge}$ is designed with the utilization of two MLPs $\phi_{s}$ and $\phi_{o}$ sharing the same architecture but with different sets of weights. These MLPs receive several inputs: the initial input node feature vectors $V_{O}^{(0)}$, the enriching object $\hat{o}$, and hidden vectors $V_{O}^{(L)}$ extracted from the final layer of $G_{sg,obj}$. The MLPs $\phi_{s}$ and $\phi_{o}$ are specifically designed to capture information from the objects and subjects, respectively, within the context of directed inter-object relationships. This process is clearly illustrated in Eq. 8 and further elucidated in the supplementary material in Fig. 2.
@@
In the Enriching Edge Detector $G_{sg,edge}$, each node vector is transformed into another space to determine the location of the new edges within the enriched graph. These transformed object vectors are then compared using the dot product to detect the enriching edges. The similarity of node features in the new space indicates a higher probability of an edge existing between the two corresponding nodes, as reflected in the entries of $\hat{M}$, similar to an adjacency matrix. Consequently, nodes denoted as $o_{e}$ are selected that form edges $\hat{e}=(o_{e},.,\hat{o})$ or $\hat{e}=(\hat{o},.,o_{e})$, where the notation ”.” signifies a placeholder for the unknown predicate. This selection process for nodes forming enriching edges by considering the similarity of node features in the transformed space is summarized as:
@@
After the determination of edges between the enriching node $\hat{o}$ and the pre-existing objects is completed, the newly formed graph $\tilde{g}=\{g,\hat{o},\hat{e}\}$ is then passed to $G_{sg,pred}$. Notably, $G_{sg,pred}$ maintains the same architecture as $G_{sg,obj}$ but operates with an independent set of parameters:
@@
Finally, a new graph $\hat{g}$ is assembled by combining the input graph $g$, enriching object $\hat{o}$, the edges $\hat{e}$, and the relationships $\hat{r}$. This comprehensive graph serves as input for the following enrichment iterations and represents the enriched scene with all the appended elements and their associated attributes.
III-C2 SG Critic.
Scene Graph Critic $D_{sg}$, as depicted in Fig. -1, indeed consists of a pair of discriminators trained jointly with the rest of the network. Some additional details regarding the architecture of GCNs are provided in the supplementary materials.
@@
In the global discriminator denoted as $D_{sg,glob}$, a GCN is employed to transform the input into another graph. The extracted features in nodes and edges of this graph are utilized to differentiate the source of input data. Given that our model handles graphs with varying numbers of nodes and edges, an aggregation method such as averaging needs to be utilized at this stage. The output vectors for nodes and edges are separately aggregated and then input into additional neural network layers for further processing and discrimination.
@@
Our global discriminator is deeper compared to the local discriminator represented as $D_{sg,loc}$ by incorporating two additional GConv layers. The local discriminator exclusively receives a subgraph comprising the enriching nodes, their immediate neighbors, and corresponding edges. The outputs of two discriminators are concatenated and subsequently passed through additional layers. This process ultimately culminates in the computation of a score. This score reflects the discrimination of an enriched graph from an original one in our critic’s assessment.
III-D Stage 2: Image Generation
Our base image generator employs a pre-trained image synthesizer [3]. Precisely, the process entails feeding scene graphs to a GCN for object features to be extracted. These extracted features are then utilized to form a scene layout for image synthesis with some convolutional layers. Simultaneously, image discriminators are jointly trained to ensure that the objects within the generated image are recognizable and image patches appear realistic, enhancing the overall quality and coherence of the generated images.
@@ III-D Stage 2: Image Generation
Both the GT scene graph $g_{GT}$ corresponding to the input graph $g$ in addition to the resulting enriched scene graph $\hat{g}$ are provided as input to the image generator $G_{im}$ to produce images $I$ and $\hat{I}$ (Eq. 11). These images further contribute to the overall enrichment objective function.
$I=G_{im}(g_{GT});\hat{I}=G_{im}(\hat{g}).$ (11)
III-E Stage 3: Visual Evaluation & Alignment
III-E1 Visual Scene Characterizer.
@@ III-E Stage 3: Visual Evaluation & Alignment
We need to ensure that the enriched output image, with its additional content, faithfully reflects the essential characteristics outlined in the GT description. Thus, a pre-trained scene classifier $S_{cf}$ [96] receives the generated enriched image $\hat{I}$ to evaluate the mentioned criteria. The hidden scene features for both images $I$ and $\hat{I}$ are extracted (Eq. 12). These extracted features are afterward integrated into the objective function, thus playing a crucial role in the end-to-end training of all components within the framework.
$h_{I}=S_{cf}(I);h_{\hat{I}}=S_{cf}(\hat{I}).$ (12)
III-E2 Image-Text Aligner.
@@
The Image-Text Aligner, represented as $M_{im\_sg}$, leverages pre-trained CLIP encoders [97]. It aims to ensure that the enriched generated images accurately reflect the essential scene elements outlined in the input description. This multimodal network operates by accepting both the input graph $g$ and the enriched image $\hat{I}$. It extracts features denoted as $F_{g}$ from the graph and $F_{I}$ from the image:
$F_{g}=M_{im\_sg}(g);F_{I}=M_{im\_sg}(\hat{I}).$ (13)
@@
III-F Objective Function
To train the model end-to-end, one random object is selected from each GT graph $g_{GT}$ to be eliminated along with its connected edges. $G_{sg,obj}$ aims to predict the object category of this removed node, which leads to our Objs Loss,
@@ III-F Objective Function
Based on $\hat{M}$, one of the edges could be selected, and then its predicate is predicted by $G_{sg,pred}$. In the GT, the predicate for this edge may or may not be provided, leading to the formulation of two supplementary loss terms computed via cross-entropy as Available Preds Loss and Not Avail Preds Loss:
$\mathcal{L}_{pred,avail}=\frac{1}{N}\sum_{i=1}^{N}CE(\hat{r}^{(i)},r_{GT}^{(i)}),$ (16)
$\mathcal{L}_{pred,not\_avail}=\frac{1}{N}\sum_{i=1}^{N}CE(\hat{r}^{(i)},none\_pred),$ (17)
@@
$\begin{split}\mathcal{L}=w_{0}\mathcal{L}_{obj}+w_{1}\mathcal{L}_{edg}+w_{2}\mathcal{L}_{GAN}+w_{3}\mathcal{L}_{pred,avail}\\ +w_{4}\mathcal{L}_{pred,not\_avail}+w_{5}\mathcal{L}_{sc}+w_{6}\mathcal{L}_{im\_sg}\end{split}$ (22)
@@ IV Experiments
Given the absence of existing work on graphs that directly addresses the specific task we tackle, our primary focus is to demonstrate the promise of the proposed approach. We also include a comparison with ChatGPT as a Large Language Model (LLM). Our objective is to illustrate the feasibility of enriching generated content rather than surpassing state-of-the-art text/graph-to-image models, GNNs, or scene classifiers. Other end-to-end trainable modules with similar functionality can be integrated into our approach to achieve potentially enhanced performance.
@@
During the inference phase, we utilize Stable Diffusion [2], a more advanced image synthesizer compared to SG2IM [3], which is employed during the training phase. This allows for enhanced visualization of details without the computational constraints imposed during training.
IV-A Dataset
For our experiments, we employed the Visual Genome (VG) dataset [1], which comprises approximately 110k images, each accompanied by its corresponding scene graph. We adopted a preprocessing pipeline similar to that in [3] where only objects and relationships with at least 2k and 0.5 occurrences are selected. This resulted in a dataset containing 178 object categories and 45 predicate categories. Consequently, for our training, testing, and validation splits, we retained approximately 62.5k, 5.5k, and 5k data points, respectively.
IV-B ChatGPT Comparison Experiments
@@
In case the input is represented as raw text rather than modeling as scene graphs, LLMs may be employed to expand the scene descriptions [98, 99, 100, 101, 102, 103, 104]. However, LLMs cannot efficiently and effectively perform the GCE task. First, simultaneous training of LLMs and image generators demands substantial training data and entails significant computational expenses. Secondly, despite the LLMs’ capacity to enrich the textual information, they lack the ability to comprehend the visual implications of the augmented content and measure its adequacy for generating a single real-world image. Thirdly, the direct utilization of LLMs might result in redundant content generation for image synthesis, particularly given that even the state-of-the-art generation method [2] struggles with handling exceedingly complex prompts.
@@
To illustrate the superiority of our proposed framework over LLMs, we conducted a comparison utilizing ChatGPT [99] for the content enrichment task. This is presented qualitatively in Fig. IV-B and quantitatively in Table II. The input descriptions are provided to ChatGPT, which generates enriched descriptions that are then used to synthesize images.
@@
We limited the word count for ChatGPT output as originally it generates very long texts unsuitable for inputs to SOTA image generators. These three different styles of prompts were used:
  • ChatGPT Direct: Please complete this given text so that the final text contains between 15 to 25 words: ”input_description
  • ChatGPT Scene: Please complete this given text describing a scene so that the final text contains between 15 to 25 words: ”input_description
  • ChatGPT Object: Please complete this given text describing a scene with more related objects so that the final text contains between 15 to 25 words: ”input_description
@@
IV-C Qualitative Results
Fig. IV-B showcases input descriptions and their enrichments alongside their corresponding synthesized images employing ChatGPT Direct and our proposed model. These input scene graphs are selected from the VG test split. The input scene graphs, along with the enriched scene graphs generated by our model, are represented as textual descriptions in Fig. IV-B (a) and (c).
@@ IV-C Qualitative Results
To ensure there is no overlap in VG between the data used for training different stages of our model, Stage1: Image Generation is pre-trained on the same train split as ours. Moreover, in Visual Scene Characterizer, we employ a pre-trained residual CNN scene classifier [96]. During network training, the gradients backpropagate through Stage 3: Visual Evaluation & Alignment and Stage2: Image Generation; however, their parameters are not updated, as our objective function is specifically designed to optimize a different problem which is joint SG/image enrichment.
IV-D Quantitative Results
@@
Objs Acc 19.04 17.31 15.09 16.85
Avail Preds Acc 41.95 39.90 40.25 31.38
Not Avail Preds Acc - - 3.02 -
Avail Edges Acc 75.00 75.64 76.08 66.89
Not Avail Edges Acc 99.58 99.70 99.54 99.54
Scene Class Acc 20.33 20.84 19.18 25.33
TABLE I: Quantitative Results, including our defined measures for the Generated Contents Enrichment task on scene graphs, is shown for selected hyperparameter configurations. For more detailed information, please refer to the supplementary materials.
@@
IV-D1 Inpainting Inspired Metrics
@@
As so many terms are encompassed within our loss function, the contribution of each term plays an essential role in the overall enrichment task. For this reason, we conducted comprehensive hyperparameter tuning. This includes the selected hyperparameters corresponding to the models in Table I, providing our quantitative evaluation methods for some of the chosen instances. SG Enrich denotes our proposed method with optimized hyperparameters. Deep G refers to a version utilizing a deeper generator architecture. Best NoD represents the optimal scenario without the introduced discriminator, while 128 BS signifies the use of a significantly larger mini-batch size. Detailed results and discussions can be found in the supplementary materials.
We have included Table I to demonstrate that our proposed method is capable of content enrichment in a structurally reasonable manner, as opposed to merely adding random content. It is important to emphasize that the enrichment process is inherently ill-posed since any appropriate enrichment is correct, and the final result is not deterministic. Therefore, the relatively low prediction accuracy does not provide a comprehensive illustration of the real performance of our introduced approach.
The metric Objs Acc represents the accuracy of our prediction among approximately 178 object categories. While it may appear relatively low, it is important to note that this outcome arises from the nature of our task. We randomly eliminate an object from a scene and then attempt to predict it based on the semantics and structure of the scene. Moreover, the GT scene graph is not the only reasonable answer. Thus, different objects, other than the GT, may be appended to form visually plausible enriched images while maintaining the reasonable structure and coherence of the entire scene.
@@
Avail Edges Acc measures the accuracy of predicting the presence of a directed edge between two nodes connected as subject and object in a relationship within the GT graph. Similarly, Not Avail Edges Acc represents the accuracy for edges whose existence is not provided in the GT graph. It is worth mentioning that the absence of an edge in the GT graph does not necessarily imply that predicting such an edge would lead to a description of a different scene. Therefore, any comparisons between the results should be made cautiously, as the problem inherently allows for multiple valid solutions.
In cases where the Enriching Edge Detector fails to accurately predict the existence of an enriching edge compared to the GT, we also penalize the Enriching Predicate GCN. This is achieved using the appropriate loss term and the introduced $none\_pred$ as the GT predicate category. The accuracy of this prediction is denoted as Not Avail Preds Acc, which is included in Table I only if the edge prediction fails for at least one enriching object, as per the GT graph. However, assume the Enriching Edge Detector successfully performs its task based on the GT. Then, we consider the prediction of the Enriching Predicate GCN as Avail Preds Acc. This metric assesses the accuracy of predicting the predicates among approximately 45 different relationship categories.
Scene Class Acc demonstrates whether the scene categories for a pair of images generated from GT and enriched scene graph are the same. During training, images have a lower resolution, and the pre-trained scene classifier does not have a high classification accuracy on its own. Thus, we do not anticipate a high matching accuracy for Scene Class Acc. Though the GT scene category is not unique, our goal is to enrich a scene by appending only one object and its relationships to finally synthesize images with the same essential visual scene characteristics as the GT graph.
IV-D2 Generated Image Quality Metrics
Metric
(Reference Distribution)

### v3

@@ 3.3.1 SG Enricher
Our Scene Graph Enricher $G_{sg}$ processes the input graph $g$ as $\hat{g}=G_{sg}(g)$. This operation yields an enriched graph $\hat{g}$. Afterward, $\hat{g}$ is reintroduced into the $G_{sg}$ process, initiating an iterative procedure for further enrichment. This recursive approach allows for the scene’s gradual improvement and refinement.
$G_{sg,obj}$ is modeled with the proposed GCN. Additional layers and skip connections are introduced to deepen the GCN, avoiding issues like vanishing gradients. $G_{sg,obj}$ results in the creation of an enriching object denoted as $\hat{o}$, where $\hat{o}=G_{sg,obj}(g)$. This newly generated enriching object is then added to the original graph $g$. The augmented graph, consisting of both the original objects and the newly created object, continues to serve as the input for the succeeding steps of the enrichment process.
$G_{sg,edge}$ is designed with the utilization of two MLPs $\phi_{s}$ and $\phi_{o}$ sharing the same architecture but with different sets of weights. These MLPs receive several inputs: the initial input node feature vectors $V_{O}^{(0)}$, the enriching object $\hat{o}$, and hidden vectors $V_{O}^{(L)}$ extracted from the final layer of $G_{sg,obj}$. The MLPs $\phi_{s}$ and $\phi_{o}$ are specifically designed to capture information from the objects and subjects, respectively, within the context of directed inter-object relationships. This process is clearly illustrated in Eq. 4 and further elucidated in the supplementary material.
@@
In the Enriching Edge Detector $G_{sg,edge}$, each node vector is transformed into another space to determine the location of the new edges within the enriched graph. These transformed object vectors are then compared using the dot product to detect the enriching edges. The similarity of node features in the new space indicates a higher probability of an edge existing between the two corresponding nodes, as reflected in the entries of $\hat{M}$, similar to an adjacency matrix. Consequently, given the relevant entries of $\hat{M}$ related to $\hat{o}$, we select nodes such as $o_{e}$ that pre-exist in $g$ and form edges $\hat{e}=(o_{e},.,\hat{o})$ or $\hat{e}=(\hat{o},.,o_{e})$, where the notation “.” represents a placeholder for an unknown predicate. This selection process for nodes forming enriching edges by considering the similarity of node features in the transformed space is summarized as:
@@
After the determination of edges between the enriching node $\hat{o}$ and the pre-existing objects $o_{e}$ is completed, the newly formed graph $\tilde{g}=\{g,\hat{o},\hat{e}\}$ is then passed to $G_{sg,pred}$. Notably, $G_{sg,pred}$ maintains the same architecture as $G_{sg,obj}$ but operates with an independent set of parameters as
@@
Finally, the new graph $\hat{g}$ is assembled by combining the input graph $g$, enriching object $\hat{o}$, the edges $\hat{e}$, and the relationships $\hat{r}$. This comprehensive graph serves as input for the following enrichment iterations and represents the enriched scene with all the appended elements and their associated attributes.
3.3.2 SG Critic
@@
Scene Graph Critic $D_{sg}$, as depicted in Fig. 2, indeed consists of a pair of discriminators trained jointly with the rest of the network. Some additional details regarding the architecture of GCNs are provided in the supplementary materials. In the global discriminator denoted as $D_{sg,glob}$, a GCN is employed to transform the input into another graph. The extracted features in nodes and edges of this graph are utilized to differentiate the source of input data. Given that our model handles graphs with varying numbers of nodes and edges, an aggregation method such as averaging needs to be utilized at this stage. The output vectors for nodes and edges are separately aggregated and then input into additional neural network layers for further processing and discrimination.
@@
Our global discriminator is deeper compared to the local discriminator represented as $D_{sg,loc}$ by incorporating two additional GConv layers. The local discriminator exclusively receives a subgraph comprising the enriching nodes $\hat{o}$, their immediate neighbors, and the corresponding edges. The outputs of two discriminators are concatenated and subsequently passed through additional layers. This process ultimately culminates in the computation of a score. This score reflects the discrimination of an enriched graph from an original one in our critic’s assessment.
3.4 Stage 2: Image Generation
@@
Our base image generator employs a pre-trained image synthesizer (Johnson et al., 2018). The process entails feeding scene graphs to a GCN for object features to be extracted. These extracted features are then utilized to form a scene layout for image synthesis with some convolutional layers. Both the GT scene graph $g_{GT}$ corresponding to the input graph $g$ in addition to the resulting enriched scene graph $\hat{g}$ are provided as input to the image generator $G_{im}$ to produce images $I$ and $\hat{I}$ (Eq. 7). These images further contribute to the overall enrichment objective function.
$I=G_{im}(g_{GT}),\;\;\;\hat{I}=G_{im}(\hat{g}).$ (7)
3.5 Stage 3: Visual Evaluation & Alignment
3.5.1 Visual Scene Characterizer
@@
We need to ensure that the enriched output image, with its additional content, faithfully reflects the essential characteristics outlined in the GT description. Thus, a pre-trained scene classifier $S_{cf}$ (Zhou et al., 2017) receives the generated enriched image $\hat{I}$ to evaluate the mentioned criteria. The hidden scene features for both images $I$ and $\hat{I}$ are extracted (Eq. 8). These extracted features are afterward integrated into the objective function, thus playing a crucial role in the end-to-end training of all components within the framework.
$h_{I}=S_{cf}(I),\;\;\;h_{\hat{I}}=S_{cf}(\hat{I}).$ (8)
3.5.2 Image-Text Aligner
@@
The Image-Text Aligner, represented as $M_{im\_sg}$, leverages pre-trained CLIP encoders (Radford et al., 2021). It aims to ensure that the enriched generated images accurately reflect the essential scene elements outlined in the input description. This multimodal network operates by accepting both the input graph $g$ and the enriched image $\hat{I}$. It extracts features denoted as $F_{g}$ from the graph and $F_{I}$ from the image:
$F_{g}=M_{im\_sg}(g),\;\;\;F_{I}=M_{im\_sg}(\hat{I}).$ (9)
@@
3.6 Objective Function
To enable end-to-end training, we randomly eliminate one object from each GT scene graph $g_{GT}$, along with its associated edges. The modified graph is then used as the input scene graph $g$. The goal of $G_{sg,obj}$ is to predict the object category of this removed node, which leads to our Obj. Loss,
@@ 3.6 Objective Function
Based on $\hat{M}$, one of the edges could be selected, and then its predicate is predicted by $G_{sg,pred}$. In the GT, the predicate for this edge may or may not be provided, leading to the formulation of two supplementary loss terms computed via cross-entropy as Available Preds Loss and Not Avail. Pred. Loss:
$\mathcal{L}_{pr,a}=\frac{1}{N}\sum_{i=1}^{N}CE(\hat{r}^{(i)},r_{GT}^{(i)}),\;\;\;\mathcal{L}_{pr,na}=\frac{1}{N}\sum_{i=1}^{N}CE(\hat{r}^{(i)},none\_pred),$ (12)
@@
$\begin{split}\mathcal{L}=w_{0}\mathcal{L}_{obj}+w_{1}\mathcal{L}_{edge}+w_{2}\mathcal{L}_{GAN}+w_{3}\mathcal{L}_{pr,a}+w_{4}\mathcal{L}_{pr,na}+w_{5}\mathcal{L}_{sc}+w_{6}\mathcal{L}_{im\_sg}\end{split}$ (17)
@@ 4 Experiments
Given the absence of existing work on graphs that directly addresses the specific task we tackle, our primary focus is to demonstrate the promise of the proposed approach. We also include a comparison with ChatGPT as a Large Language Model (LLM). Our objective is to illustrate the feasibility of enriching generated content rather than surpassing SOTA text/graph-to-image models, GNNs, or scene classifiers. Other end-to-end trainable modules with similar functionality can be integrated into our approach to achieve potentially enhanced performance.
@@
During the inference phase, we utilize Stable Diffusion (Rombach et al., 2022), a more advanced image synthesizer compared to SG2IM (Johnson et al., 2018), which is employed during the training phase. This allows for enhanced visualization of details without the computational constraints imposed during training.
4.1 Dataset
We utilized the Visual Genome (VG) dataset (Krishna et al., 2017), which consists of 110k real-world images with their scene graphs. We employed a preprocessing similar to Johnson et al. (2018). This resulted in 178 object and 45 predicate categories in approximately 62.5k, 5.5k, and 5k samples for training, testing, and validation, respectively. VG offers rich and detailed annotations, comprising 3.8M objects and 2.3M relationships, which capture a diverse range of environments and object configurations
The COCO-Stuff dataset (Caesar et al., 2018) with only 25k training images, is often used alongside VG (Johnson et al., 2018; Yang et al., 2022c; Herzig et al., 2020). Like VG, COCO-Stuff selected its images from the MS COCO dataset (Lin et al., 2014). However, COCO-Stuff lacks scene graphs and only provides bounding boxes. Thus, scene graphs are synthesized based on the 2D coordinates of objects, with only six automatically generated relationships: left of, right of, above, below, inside, and surrounding.
Due to the synthetic nature and limited relational scope of COCO-Stuff, we believe it is unsuitable for demonstrating the full potential of our model’s enrichment capabilities, as the generated graphs do not accurately reflect real-world descriptions. To the best of our knowledge, no other non-synthetic scene graph datasets are available for our purposes. In contrast, the VG dataset’s detailed and realistic annotations enable our model to generalize effectively, generating a wide variety of scenes with complex and diverse object relationships.
4.2 ChatGPT Comparison Experiments
@@
In case the input is represented as raw text rather than modeling as scene graphs, LLMs may be employed to expand the scene descriptions (Brown et al., 2020; OpenAI, 2023; Touvron et al., 2023; Scao et al., 2022; Ouyang et al., 2022; Anil et al., 2023; Driess et al., 2023). However, LLMs cannot efficiently and effectively perform the GCE task. First, simultaneous training of LLMs and image generators demands substantial training data and entails significant computational expenses. Secondly, despite the LLMs’ capacity to enrich the textual information, they lack the ability to comprehend the visual implications of the augmented content and measure its adequacy for generating a single real-world image. Thirdly, the direct utilization of LLMs might result in redundant content generation for image synthesis, particularly given that even the SOTA generation method (Rombach et al., 2022) struggles with handling exceedingly complex prompts.
@@
To illustrate the superiority of our proposed framework over LLMs, we conducted a comparison utilizing ChatGPT (OpenAI, 2023) for the content enrichment task. This is presented qualitatively in Fig. 3 and quantitatively in Table 2. The input descriptions are provided to ChatGPT, which generates enriched descriptions that are then used to synthesize images.
@@
We limited the word count for ChatGPT output as originally it generates very long texts unsuitable for inputs to SOTA image generators. These three different styles of prompts were used: (i) ChatGPT Direct: Please complete this given text so that the final text contains between 15 to 25 words: "input_description". (ii) ChatGPT Scene: Please complete this given text describing a scene so that the final text contains between 15 to 25 words: "input_description". (iii) ChatGPT Object: Please complete this given text describing a scene with more related objects so that the final text contains between 15 to 25 words: "input_description".
4.3 Qualitative Results
@@
Fig. 3 showcases input descriptions and their enrichments alongside their corresponding synthesized images employing ChatGPT Direct and our proposed model. These input scene graphs are selected from the VG test split. The input scene graphs, along with the enriched scene graphs generated by our model, are represented as textual descriptions in Fig. 3 (a) and (c).
@@
To ensure there is no overlap in VG between the data used for training different stages of our model, Stage1: Image Generation is pre-trained on the same train split as ours. Moreover, in Visual Scene Characterizer, we employ a pre-trained residual CNN scene classifier (Zhou et al., 2017). During network training, the gradients backpropagate through Stage 3: Visual Evaluation & Alignment and Stage2: Image Generation; however, their parameters are not updated, as our objective function is specifically designed to optimize a different problem which is joint SG/image enrichment.
4.4 Quantitative Results
Table 1: Our defined measures for prediction accuracy are presented across selected hyperparameter configurations. The supplementary materials provide details on the models. The enrichment task is inherently ill-posed, allowing for multiple correct solutions. However, our method enriches content in a structurally reasonable manner, avoiding random generation.
@@ 4.4 Quantitative Results
Obj. Acc. 19.04 17.31 15.09 16.85
Avail. Pred. Acc. 41.95 39.90 40.25 31.38
Not Avail. Pred. Acc. - - 3.02 -
Avail. Edge Acc. 75.00 75.64 76.08 66.89
Not Avail. Edge Acc. 99.58 99.70 99.54 99.54
Scene Class. Acc. 20.33 20.84 19.18 25.33
@@
4.4.1 Inpainting Inspired Metrics
@@
As many terms are encompassed within our loss function, the contribution of each term plays an essential role in the overall enrichment task. For this reason, we conducted comprehensive hyperparameter tuning. This includes the selected hyperparameters corresponding to the models in Table 1, providing our quantitative evaluation methods for some of the chosen instances. SG Enrich denotes our proposed method with optimized hyperparameters. Deep G refers to a version utilizing a deeper generator architecture. Best NoD represents the optimal scenario without the introduced discriminator, while 128 BS signifies the use of a significantly larger mini-batch size. Detailed results and discussions can be found in the supplementary materials. Table 1 demonstrates that our proposed method is capable of content enrichment in a structurally reasonable manner, as opposed to merely adding random content. It is important to emphasize that the enrichment process is inherently ill-posed since any appropriate enrichment is correct, and the final result is not deterministic. Therefore, the relatively low prediction accuracy does not provide a comprehensive illustration of the real performance of our introduced approach.
The metric Obj. Acc. represents the accuracy of our prediction among approximately 178 object categories. While it may appear relatively low, it is important to note that this outcome arises from the nature of our task. We randomly eliminate an object from a scene and then attempt to predict it based on the semantics and structure of the scene. Moreover, the GT scene graph is not the only reasonable answer. Thus, different objects, other than the GT, may be appended to form visually plausible enriched images while maintaining the reasonable structure and coherence of the entire scene.
@@
Avail. Edge Acc. measures the accuracy of predicting the presence of a directed edge between two nodes connected as subject and object in a relationship within the GT graph. Similarly, Not Avail. Edge Acc. represents the accuracy for edges whose existence is not provided in the GT graph. It is worth mentioning that the absence of an edge in the GT graph does not necessarily imply that predicting such an edge would lead to a description of a different scene. Therefore, any comparisons between the results should be made cautiously, as the problem inherently allows for multiple valid solutions.
In cases where the Enriching Edge Detector fails to accurately predict the existence of an enriching edge compared to the GT, we also penalize the Enriching Predicate GCN. This is achieved using the appropriate loss term and the introduced $none\_pred$ as the GT predicate category. The accuracy of this prediction is denoted as Not Avail. Pred. Acc., which is included in Table 1 only if the edge prediction fails for at least one enriching object, as per the GT graph. However, assume the Enriching Edge Detector successfully performs its task based on the GT. Then, we consider the prediction of the Enriching Predicate GCN as Avail. Pred. Acc.. This metric assesses the accuracy of predicting the predicates among approximately 45 different relationship categories.
Scene Class. Acc. demonstrates whether the scene categories for a pair of images generated from GT and enriched scene graph are the same. During training, images have a lower resolution, and the pre-trained scene classifier does not have a high classification accuracy on its own. Thus, we do not anticipate a high matching accuracy for Scene Class. Acc.. Though the GT scene category is not unique, our goal is to enrich a scene by appending only one object and its relationships to finally synthesize images with the same essential visual scene characteristics as the GT graph.
4.4.2 Generated Image Quality Metrics
Table 2: Inception Scores (IS) (Salimans et al., 2016) and Fréchet Inception Distances (FID) (Heusel et al., 2017) are computed to compare real-world VG dataset and sets of images generated from GT scene descriptions (GT Synthesized), Simple descriptions, and Enriched versions utilizing Our model and ChatGPT with three different prompt styles.
Metric [Reference Distribution]